In a right triangle, one leg is twice the other leg, and the area is 9. How long is the hypotenuse?
Answer by ankor@dixie-net.com(22740)

In a right triangle, one leg is twice the other leg, and the area is 9. How long is the hypotenuse?
Let x = one side, the other side = 2x
:
Area of a triangle = (1/2)*b*h; in a right triangle, the base and the height
are the same as the sides:
:
(1/2)*x*2x = 9
(1/2)*2x^2 = 9
x^2 = 9
x = 3
:
hypotenuse = SqRt(3^2 + 6^2) ~ 6.71